OpenStudy (anonymous):

If I buy 2 apples and 6 oranges, I will spend $5.60. However, if I buy 6 apples and 2 oranges, I will spend $4.80. How much does one orange cost?

OpenStudy (unklerhaukus):

\[2a+6o=$5.60\]\[6a+2o=$4.80\]

OpenStudy (anonymous):

how much does each orange cost

OpenStudy (unklerhaukus):

hmmm you might have to solve the simultaneous equations

OpenStudy (anonymous):

can u help me please:(

OpenStudy (unklerhaukus):

sure

OpenStudy (anonymous):

0k thank u sweetheart

OpenStudy (unklerhaukus):

the first step can be to rearrange this equation\[2a+6o=$5.60\] so that a (a for apples) is only on one side to do this take away 6o from both sides and then dived both sides by two

OpenStudy (anonymous):

i dont understand sorry

OpenStudy (unklerhaukus):

\[2a=$5.60-6o\] \[a=\frac{$5.60-6o}2\]\[ a =$2.80-3o\]

OpenStudy (unklerhaukus):

does the first step make sense now?

OpenStudy (anonymous):

no

OpenStudy (unklerhaukus):

what is the confusing part (s)

OpenStudy (anonymous):

im a very slow learner im sorry for wasting your good time sir

OpenStudy (unklerhaukus):

i am happy to go at your pace

OpenStudy (anonymous):

no help me

OpenStudy (unklerhaukus):

what is the confusing part

OpenStudy (anonymous):

im jewish bye